OpenStudy (anonymous):

If a region on the coordinate plane is enclosed by the functions y = 0, x = 1, x =3, and y = (4/x), and is to be rotated about the line y = -1, what is the volume of this region? I know that this is going to look like a washer (or hollow ring), but I am having some difficulty coming up with the equations for the inner and outer radius I know that the outer function is y = 4/x, inner function is y = 0....

OpenStudy (anonymous):

@TuringTest Would you mind helping me out with this one? I think inner radius = -1 - (0) I think outer radius = -1 - (4/x)

OpenStudy (anonymous):